Back to the NIKA2 main page

Daily Reports

Mar 1st, Tuesday

22:00-03:00; Frédéric (on-site) & Hélène (remote): [IM: We start at 22:00, after maintenance and the technical time slot T06-21 that was dedicated to measurements for the telescope efficiencies. Initial NIKA2 tuning looks fine with 29/7/50 bad KIDs for A1/A2/A3, respectively. However, initial pointing and focus yields a strange effect only for 1mm data: 4 images of the pointing source in a cross-like pattern. This was not visible for 2mm data. We repeat the pointing to another source and get the same effect. Then we try to re-start everything: put mirrors back to EMIR setup and then back to NIKA2 setup and restart DAQ and retune KIDs. After these steps the pattern at 1mm data goes away. We have good beam size and shape (~17.7" @ 2mm and ~12" @ 1mm) as well as atmospheric fluctuations are very low (~0.5 Jy/beam). After a calib_1scan [20220301s246] and a beam map [20220301s247] on the quasar 0316+413, we start with observations for the IMEGIN LP (160-16).] Earlier that evening the staff performed some tests to characterize the efficiency of the telescope. The AoD started tuning NIKA2 around 22:05. After a series of calibrations, we realized that the 1mm beam was shaped as a quadrupod (two images in azimut, two images in elevation), although the 2mm beam was perfectly normal. The operator went to change the configuration several times and the AOD retuned NIKA2, until the problem disappeared. A beammap was launched on 0316+413 at 23:40 after a calib_1scan. We started observing project 160-16 at 00:30. We stayed the whole shift on NGC4321 (@60"/s; 9 complete scans in total; 101 minutes on source), but were blocked in the middle of scan 20220302s13 without noticing, and for an unknown reason. An alarm was triggered, but nobody realized its cause. The elevation was less than 70 degrees, so the scan speed should not have been a problem. We calibrated on 1226+033 (pointing and focus) at 00:10 and 02:03. The focus changed from -0.35 mm to -0.60 mm in 3 hours. The opacity was very stable at 0.10 to 0.11 for the whole shift, and the sky rms at 2mm was between 0.5 and 1 Jy/beam. Too bad a lot of time was lost in such a fantastic weather!

Here is an example with the strange effect at 1mm data. Both plots correspond to the same scan (20220301s241), one for Ar1 and the other for Ar2:

attachment:20220301s241_Ar1.png attachment:20220301s241_Ar2.png

RZ, 2022.03.02 9:49

1. the beam of ar2 is not affected => problem is only of the 1mm arrays => not the mirrors

2. the source is shifted only in the scanning direction: in az in the focus maps, in az&el in pointings => synchro problems of the 1mm arrays

the angular distance btw. the shifted sources + the scanning velocity allow to calculate the synchro. time error, from the pointings: ~25"/(2*25"/s)=0.5s

Mar 2nd, Wednesday

03:00-05:00; Miren: Took over at 2h UT with tau255 ~ 0.125. Starting with pointing. Less than +/- 1 Jy/beam fluctuations and 17.4’’ FWHM beam @2mm. Project 199-16: one repetition on PSZ2G081 (40.5/44) and a repetition on PSZ2G046 (17.5/20). Pointing and focus, followed by a skydip (20220302s32). Sky fluctuations +/- 5 Jy/beam @ 2mm and tau225 ~ 0.15, 17.75’’ FWHM beam @2mm. Another repetition on PSZ2G046 (18.5/20). Ioannis takes over at 4h15 UT.

05:00-10:30; Ioannis: EMIR polarization observations with the POLAMI LP (087-21)

10:30-15:00; Philippe: Clear sky and fairly good weather conditions with tau225 ~ 0.2-0.3. Stable focus -1.2 to -1.0mm. Started with NIKA2-Pol around 11:00, following an X-POL (Polami) session. With NIKA2-Pol, took two calib_1scans on 3C345 (s118 & s133), two calib_1scans on Mars (s120 and s126), 1 sequence of 3 maps on Serp-SMM4 and 2 sequences of 3 maps on DR21OH for the B-FUN project (015-17). Also observed a polarized skydip (s145).

15:00-21:00; Aurore: Αverage weather conditions with some clouds --> switched to EMIR project 116-21, receiver E0. Made in total 24 repetitions from ~ 15h-21h. Water vapor around 5 mm during the whole shift, increasing towards the end. PaKo's green screen froze once (but commands were still taken into account). Restarting pako solved the problem. Pointings were good, focus also fine, although the atmosphere was a little unstable.

21:00-03:00; Frédéric (on-site), with help from Hélène (remote) during the first hour: We stayed on project 170-21 with EMIR for the whole shift. We did 8 repetitions on 8728-high and 9 repetitions on 3670-high. We did one pointing every hour. The focus was very stable over night. The opacity@225 was slowly rising from 0.4 and 0.7 during the shift, with a few spikes above 0.9 after 00:30.

Mar 3rd, Thursday

03:00-09:00; Miren: [IM: EMIR backup observations of project 170-21]

09:00-15:00; Philippe: [IM: EMIR backup observations until 13:00. Then stopped due to bad weather: fog, snow] Cloudy, with Tau225 ~ 0.6-0.8. EMIR backup observations of projects 170-21 and 174-21. Had to stop observing with EMIR after 13:00 as pointing was not possible even on strong sources...

15:00-21:00; Ioannis: Stopped due to bad weather.

21:00-03:00; Frédéric: [IM: Stopped due to bad weather.]

Mar 4th, Friday

IM: Stopped all day due to bad weather, mainly fog and snow. In the night, the conditions improved a bit and the operator cleaned the subreflector in preparation of restarting. Several tries throughout the night on bright sources but the atmosphere was too unstable. No observations.

Mar 5th, Saturday

IM: Stopped all day due to bad weather. No observations.

Mar 6th, Sunday

03:00-09:00: Still stopped.

09:00-15:00; Philippe: Starting over after a long period of bad weather (snow/fog); De-icing switched off at 9:30 LT + cleaning of the dish by the operator. After a few pointing tests with EMIR, started NIKA2-Pol observations at ~11:30 LT with tau225 ~ 0.3-0.35: 1 calib_1scan on Mars (s33) + 1 calib_1scan on 1641+399=3C345 (s35), followed by a sequence of 2 maps on Serp-SMM4 (s37, s38) for the B-FUN project (015-17). Weather conditions (opacity rising above 0.5 + clouds) degrading quickly during s38. Had to switch to 3mm EMIR backup observations (project 174-21) after s38 at about 13:00 LT. Observed with EMIR for project 174-21 until ~ 14:45 LT, then had to stop as it started snowing slightly again ...

15:00-21:00; Ioannis: Tau225 between 0.3 and 0.8. Very unstable conditions, fog and light snow. No observations.

21:00-03:00; Frédéric: The opacity@225 was quite high and fluctuating, earlier the evening (0.3-0.6). At 21:45, we decided to do EMIR observations for project 170-21. We started with a calibration, focus and pointing on a bright source (0316+413). We had to reiterate the calibration 3 times, as the receiver temperature was around 200. Pointing and calibration were bad. We did several iterations under project test. Around 00:25, we finally got a decent pointing and focus. The opacity was fluctuating less (0.3-0.4). We launched two iterations on 16881-high to finish this source. We then moved to 17663-high, for 2 repetitions. We had to cancel in the middle of the second repetition, as it started snowing, at 01:10. At 02:30, snow had stopped and tau was back to 0.3. We tried another pointing on a bright source (1226+023). The conditions were stable and we launched a last iteration on 17663-high.

Mar 7th, Monday

03:00-09:00; Miren: [IM: Stopped for most of the shift due to bad weather. Only a few EMIR scans for project 170-21. Snow started towards the end of the shift.]

09:00-15:00; Philippe: No observations due to snow.

15:00-21:00; Aurore: No observations due to snow.

21:00-03:00; Frédéric: The weather was bad earlier the evening. At 00:30, the sky opacity started improving. We performed a bright source pointing and focus, which were good. We started observing project 170-21 with EMIR. We performed 2 repetitions on 17663-high. We stopped in the middle of the second one, because the README file had changed in the meantime. We do not know if we were looking at a non-refreshed README file or if someone edited it. It previously stated that 52 minutes out of 2 hours had been observed. It now said "!!NEW detected". We thus moved to 41401-low (2 repetitions). The condition improved with tau<0.3. We then moved to 58740-low (2 repetitions). tau225 was ~0.25. Pointing looked stable.

Mar 8th, Tuesday

03:00-07:30; Miren: Continued with EMIR backup observations of project 170-21 until about 06:30, when the opacity increased abruptly again and the atmosphere became very unstable.

07:30-10:30; Ioannis: Several unsuccessful attempts to restart observations with EMIR for the POLAMI large program (087-21).

DailyReportsNika2Pool010322 (last edited 2022-03-09 13:44:32 by NikaBolometer)